OpenStudy (anonymous):

What is the solution to the equation log sub 4 x – log sub 4 2 = 3

Parth (parthkohli):

\(\Rightarrow \log_4x - {1 \over 2} = 3 \) \(\Rightarrow \log _4 x = 2.5 \) \(\Rightarrow x = 4^{2.5} \)

Parth (parthkohli):

\(\Rightarrow x = 4^2 \times 4^{0.5} \) \(\Rightarrow x = 16 \times \sqrt4 \) \(\Rightarrow x = 16 \times 2 \) \(\Rightarrow x = 32\)

Parth (parthkohli):

Do you understand? @strcaitlyn3

Parth (parthkohli):

13*

OpenStudy (anonymous):

yea i understand :)

Parth (parthkohli):

Cool!

OpenStudy (anonymous):

hey 32 wasnt right. :( is said that x=32 was inccorect. i got a 92 out of a hundred.